### v1

Abstract.
[…]
Key words and phrases:
$\Gamma$-Convergence, Weakly compactly generated space, Fenchel conjugate, Convergence of subdifferentials, Separable Reduction, Rich families, Projectional skeleton
2020 Mathematics Subject Classification:
Primary 49J52; Secondary 46B26, 47N10, 90C25
Rafael Correa was partially supported by Centro de Modelamiento Matemático (CMM), ACE210010 and FB210005, BASAL funds for center of excellence and ANID-Chile grant: Fondecyt Regular 1240335, P. Pérez-Aros was partially supported by Centro de Modelamiento Matemático (CMM), ACE210010 and FB210005, BASAL funds for center of excellence and ANID-Chile grant: MATH-AMSUD 23-MATH-09 and MATH-AMSUD 23-MATH-17, ECOS-ANID ECOS320027, Fondecyt Regular 1220886, Fondecyt Regular 1240335, Fondecyt Regular 1240120 and Exploración 13220097. José Pablo Santander was partially supported by Centro de Modelamiento Matemático (CMM), ACE210010 and FB210005, BASAL funds for center of excellence and ANID-Chile grant: Fondecyt Regular 1240335
1. Introduction
In optimization theory, when studying the limit of a sequence of optimization problems, one seeks a natural notion of convergence of the data that guarantees convergence of both optimal values and minimizers. In other words, a form of variational convergence that allows one to pass to the limit in variational problems. The first concept of this kind was introduced by Wijsman in finite dimensions under the name infimal convergence [78, 79], motivated by questions in statistical decision theory [31]. In the 1970s, De Giorgi and Franzoni extended this perspective to general topological spaces under the name $\Gamma$-convergence [47], which has since become one of the central tools of the calculus of variations.
In parallel, the infinite-dimensional setting introduced new challenges due to the role of weak topologies. To address these, Mosco introduced in the late 1960s a convergence notion for lower semicontinuous convex functions—and hence for closed convex sets via their indicator functions—in order to study the stability of solutions to variational inequalities [55, 57, 56]. This notion, now known as Mosco convergence, was closely linked to the convergence of solutions of elliptic and parabolic PDEs and the associated energy functionals (see, e.g., [75, 40, 74]).
A natural line of investigation is how $\Gamma$-convergence and Mosco convergence interact with fundamental operations in convex analysis and optimization, such as Fenchel conjugation, polarity of convex sets, or subdifferentiation. The continuity of the Fenchel conjugate with respect to variational convergences was first analyzed by Wijsman [79], who proved its continuity under infimal convergence. Around the same time, Walkup and Wets [77] showed that in reflexive Banach spaces the polarity operation for closed convex cones is an isometry; as Rockafellar later noted [66, Theorem 14.4], this already implies continuity of the Fenchel transform in finite dimensions. Related connections between polarity and convex cones were further developed in [65].
The breakthrough in infinite dimensions came with Mosco’s 1971 theorem [58], establishing that Mosco convergence makes the Fenchel conjugate bicontinuous. Reflexivity itself can be characterized by this bicontinuity property [20, 18]. Mosco’s theorem thus extends Wijsman’s results and ensures continuity of conjugation also in finite dimensions. Subsequent work by Joly [48] generalized Mosco’s results within Banach spaces, while McLinden and Bergstrom [54] explored consequences in finite dimensions for convex sets, functions, and programming models. In the following decades, continuity of the Fenchel conjugate and the polar map, and generalizations of Mosco convergence to non-reflexive spaces, as well as topologies for closed (or closed and convex) sets and convex functions, were extensively studied (see, e.g., [15, 16, 17, 7, 42, 8, 35, 60, 24, 73, 23, 22, 25, 27, 21, 19, 61]), with Beer and Borwein’s result [20] showing that bicontinuity of the Fenchel conjugate is in fact equivalent to reflexivity.
A further milestone was reached by Attouch in 1977, who proved that Mosco convergence of convex functions is equivalent to the convergence of their subdifferentials in the sense of Painlevé–Kuratowski set convergence, up to a normalization constant [3]. This fundamental result, now known as Attouch’s theorem, has become a cornerstone of variational analysis. It has found numerous applications: to stability in optimization and variational problems [14, 13, 9, 10, 11, 12], to the analysis of parabolic variational inequalities [6], to numerical stability in optimization algorithms [51], and to structural results such as generalized second-order derivatives of convex functions [68, 67, 34, 33, 52, 63] and differentiability of Lipschitz set-valued mappings [39]. Attouch’s theorem also remains influential in applied areas such as nonsmooth mechanics and the study of subgradient evolution systems [5].
Today, $\Gamma$-convergence is well-recognized as a unifying framework across diverse fields of pure and applied mathematics. While its origins lie in mathematical programming, it has become an indispensable tool in the approximation of optimization problems, the study of subdifferential inclusions, consistency in stochastic programming, calculus of variations, PDEs arising as Euler–Lagrange equations of variational problems, and even extensions of probabilistic principles such as the strong law of large numbers for functions (see, e.g., [2, 69, 70, 49, 72, 38, 71, 29, 5, 4] and the references therein).
[…]
2. Outline of the Manuscript
[…]
The remaining sections are devoted to the proofs. Section 5 gathers several auxiliary results needed in the sequel. Section 6 develops the proof of Theorem 4.1, the most substantial argument of this work. The proof is first established in separable Banach spaces (Subsection 6.1) and then extended to the general case by means of separable reduction. The method relies on rich families and projectional skeletons: Subsection 6.2 introduces the necessary concepts, while Subsection 6.4 develops the construction of rich families that remove the separability assumption. Although it is used for technical purposes in this paper, it should be noted that Corollary 6.9 is of independent interest, as it goes in line with the characterization of Asplund spaces given in [37]. Finally, Subsection 6.5 completes the argument in weakly compactly generated Banach spaces. Section 7 presents the short proof of Corollary 4.6, and Section 8 concludes with the proof of Theorem 4.7.
3. Mathematical Preliminaries
3.1. Notations and basic definitions
In what follows, $X$ will denote a Banach space with norm $\left\lVert\cdot\right\rVert$. Its dual space will be denoted by $X^{\ast}$, and $\left\lVert\cdot\right\rVert_{\ast}$ stands for the norm in $X^{\ast}$. The open ball of radius $\varepsilon>0$ centered at $\bar{x}$ is denoted by $\mathbb{B}_{\varepsilon}(\bar{x})$. For a subset $A\subset X$, $\operatorname{\overline{sp}}(A)$ will denote the closed linear span generated by $A$.
We denote by $w$ the weak topology of $X$, and by $w^{\ast}$ the weak topology of $X^{\ast}$. If a sequence $(x_{n})_{n\in\mathbb{N}}$ in $X$ converges strongly to $x\in X$, we write $x_{n}\to x$. Weak convergence in $X$ is denoted by $x_{n}\rightharpoonup x$. Similarly, if $(x_{n}^{\ast})_{n\in\mathbb{N}}\subset X^{\ast}$ converges strongly to $x^{\ast}\in X^{\ast}$, we write $x_{n}^{\ast}\to x^{\ast}$, whereas $x_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}$ denotes weak convergence.
The Mackey topology $\mu^{\ast}:=\mu(X^{\ast},X)$, defined on $X^{\ast}$, is the topology of uniform convergence on the family of all convex, symmetric, and weakly compact subsets of $X$. For a net $(x_{\alpha}^{\ast})\subset X^{\ast}$, we write $x_{\alpha}^{\ast}\overset{\mu^{\ast}}{\to}x^{\ast}$ to denote convergence to $x^{\ast}$ in the Mackey topology.
Finally, $X$ is said to be a weakly compactly generated (WCG) space if there exists a weakly compact set $A\subset X$ such that the linear span of $A$ is dense in $X$ with respect to the norm topology. In this definition, $A$ may be assumed, without loss of generality, to be convex and symmetric (see, e.g., [43, Chapter 13]).
For a function $f\colon X\to\mathbb{R}\cup\{\infty\}$, we define its (effective) domain and epigraph as follows:
$\mathrm{dom}\,f:=\{\,x\in X:f(x)<\infty\,\},\qquad\mathrm{epi}\,f:=\{\,(x,r)\in X\times\mathbb{R}:f(x)\leq r\,\}.$
Furthermore, a function $f\colon X\to\mathbb{R}\cup\{\infty\}$ will be called proper if ${\rm dom}\kern 1.19995ptf\neq\emptyset$, and will be called convex if ${\rm epi}\kern 1.19995ptf$ is a convex set.
For a function $f\colon X\to\mathbb{R}\cup\{-\infty,\infty\}$ its Fenchel conjugate is defined by
$f^{\ast}(x^{\ast}):=\displaystyle\sup_{x\in X}\left\{\langle x^{\ast},x\rangle-f(x)\right\}.$
Furthermore, for a function $h:X^{\ast}\to\mathbb{R}\cup\{-\infty,\infty\}$, its conjugate will be understood defined on $X$ as
$h^{\ast}(x):=\displaystyle\sup_{x^{\ast}\in X}\left\{\langle x^{\ast},x\rangle-h(x^{\ast})\right\}.$
For a given set $C\subset X$ we denote by $\delta_{C}$ the indicator function of $C$, that is, the function from $X$ to $\mathbb{R}\cup\{\infty\}$ defined by
[…]
The support function of $C\subset X$, denoted by $\sigma_{C}$, is the function from $X^{\ast}\to\mathbb{R}\cup\{\infty\}$ such that $\sigma_{C}(x^{\ast}):=\sup\{\langle x^{\ast},x\rangle:x\in C\}$. Observe that $\delta_{C}^{\ast}=\sigma_{C}$.
Given a proper lower semicontinuous convex function $f\colon X\to\mathbb{R}\cup\{\infty\}$ and $\lambda>0$, we recall that the Moreau envelope of $f$ associated to $\lambda$, denoted by $f_{\lambda}$, is the following map:
$f_{\lambda}(x):=\displaystyle\inf_{y\in X}\left\{f(y)+\frac{1}{2\lambda}\left\lVert x-y\right\rVert^{2}\right\}.$
More generally, for two functions $f,g\colon X\to\mathbb{R}\cup\{\infty\}$, the inf-convolution is denoted by
$f\square g(x):=\inf\left\{f(x_{1})+g(x_{2}):x_{1}+x_{2}=x\right\}.$
For lsc proper convex function $f\colon X\to\mathbb{R}\cup\{\infty\}$ one denotes by $\partial f$ the graph of the subdifferential of $f$:
$\partial f:=\{(x,x^{*})\in X\times X^{*}:x^{*}\in\partial f(x)\}.$
where $\partial f(x)$ is the subdifferential of $f$ at $x$ and is given by:
$\partial f(x):=\{x^{*}\in X^{*}:f(x)+f^{\ast}(x^{\ast})=\langle x^{\ast},x\rangle\}.$
We will denote by $\partial f^{\ast}$ the graph of the subdifferential of $f^{\ast}$ restricted to $X^{\ast}\times X$ rather than $X^{\ast}\times X^{**}$. So $x\in\partial f^{\ast}(x^{\ast})$ if and only if $x^{\ast}\in\partial f(x)$.
3.2. Limits of sequences of functions and operators
Let $(Y,\tau)$ be a topological space, and let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of functions from $Y$ into $\mathbb{R}\cup\{\infty\}$. We define two limit functions, called the lower sequential $\Gamma$-limit and the upper sequential $\Gamma$-limit of the sequence $(f_{n})_{n\in\mathbb{N}}$ (with respect to $\tau$), as follows:
(3.1) $\begin{aligned} \Gamma(\tau)\text{-}\liminf_{n}f_{n}(x)&:=\inf\left\{\liminf_{n}f_{n}(x_{n}):x_{n}\overset{\tau}{\to}x\right\},\\ \Gamma(\tau)\text{-}\limsup_{n}f_{n}(x)&:=\inf\left\{\limsup_{n}f_{n}(x_{n}):x_{n}\overset{\tau}{\to}x\right\},\end{aligned}$
where $x_{n}\overset{\tau}{\to}x$ denotes convergence of the sequence $(x_{n})_{n\in\mathbb{N}}$ with respect to the topology $\tau$. Moreover, it is a known fact that when the topology $\tau$ is metrizable (or more general, first countable), then each infimum is indeed a minimum [29, p.30, Remark 1.26].
Furthermore, we say that the sequence $(f_{n})_{n\in\mathbb{N}}$ sequentially $\Gamma$-converges under the topology $\tau$ if
$\Gamma(\tau)\text{-}\liminf_{n}f_{n}=\Gamma(\tau)\text{-}\limsup_{n}f_{n}.$
In this case, we denote its sequential $\Gamma$-limit simply by $\Gamma(\tau)\text{-}\displaystyle\lim_{n}f_{n}$.
Since we are particularly interested in $\Gamma$-limits with respect to the norm topology, in this case, we omit the symbol with respect to the topology for the sequential $\Gamma$-limits of sequences $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$ (or defined in $X^{\ast}$ instead of $X$). That is:
$\Gamma\text{-}\liminf_{n}f_{n}$ $=\Gamma(\left\lVert\cdot\right\rVert)\text{-}\liminf_{n}f_{n},$
$\Gamma\text{-}\limsup_{n}f_{n}$ $=\Gamma(\left\lVert\cdot\right\rVert)\text{-}\limsup_{n}f_{n},$
$\Gamma\text{-}\lim_{n}f_{n}$ $=\Gamma(\left\lVert\cdot\right\rVert)\text{-}\lim_{n}f_{n}.$
Let $(S_{n})_{n\in\mathbb{N}}$ be a sequence of sets in a topological space $(Y,\tau)$. The $\tau$-sequential lower limit, denoted by $\operatorname{Li}_{\tau}S_{n}$, is the set of all $y\in Y$ for which there exists a sequence $(y_{n})_{n\in\mathbb{N}}\subset Y$ with $y_{n}\in S_{n}$ for all sufficiently large $n\in\mathbb{N}$ and $y_{n}\xrightarrow{\tau}y$. Similarly, the $\tau$-sequential upper limit, denoted by $\operatorname{Ls}_{\tau}S_{n}$, is the set of all $y\in Y$ for which there exist an increasing function $j\colon\mathbb{N}\to\mathbb{N}$ and a sequence $(y_{n})_{n\in\mathbb{N}}$ with $y_{n}\in S_{j(n)}$ such that $y_{n}\xrightarrow{\tau}y$.
We now introduce the notion of graphical convergence for set-valued mappings by identifying them with their graphs. Recall that for a set-valued mapping $A\colon Y\to Z$, its graph is defined as
${\rm gph}\kern 1.19995ptA:=\{(y,z)\in Y\times Z:z\in A(y)\}.$
Given a sequence of set-valued mappings $(A_{n})_{n\in\mathbb{N}}$ from $Y$ to $Z$, the $\tau_{Y}\times\tau_{Z}$-sequential lower limit and the $\tau_{Y}\times\tau_{Z}$-sequential upper limit are defined as the set-valued mappings
$\operatorname{Li}_{\tau_{Y}\times\tau_{Z}}A_{n}\quad\text{and}\quad\operatorname{Ls}_{\tau_{Y}\times\tau_{Z}}A_{n},$
whose graphs are the sets
$\operatorname{Li}_{\tau_{Y}\times\tau_{Z}}{\rm gph}\kern 1.19995ptA_{n}\quad\text{and}\quad\operatorname{Ls}_{\tau_{Y}\times\tau_{Z}}{\rm gph}\kern 1.19995ptA_{n},$
respectively.
Moreover, we say that a sequence of set-valued mappings $A_{n}\colon Y\rightrightarrows Z$, $n\in\mathbb{N}$, sequentially $\tau_{Y}\times\tau_{Z}$-converges to a set-valued mapping $A\colon Y\rightrightarrows Z$, and we write
(3.2) $A_{n}\xrightarrow{G(\tau_{Y}\times\tau_{Z})}A,$
if
$\operatorname{Ls}_{\tau_{Y}\times\tau_{Z}}A_{n}\subset{\rm gph}\kern 1.19995ptA\subset\operatorname{Li}_{\tau_{Y}\times\tau_{Z}}A_{n},$
which is the same as
$\operatorname{Ls}_{\tau_{Y}\times\tau_{Z}}A_{n}={\rm gph}\kern 1.19995ptA=\operatorname{Li}_{\tau_{Y}\times\tau_{Z}}A_{n}.$
4. Limits of convex functions and subdifferentials
In this section we study the relation between the epigraphical convergence of a sequence of convex functions and the $w^{\ast}$-sequential limit of its conjugate functions, in the sense of (3.1).
The first result of this section provides a characterization of the conjugate of the lower sequential epigraphical limit of a sequence of functions, as the upper sequential epigraphical limit of their corresponding conjugates with respect to the weak-topology. The proof of this result will be presented in Section 6.
Theorem 4.1.
Let $X$ be a WCG Banach space, and let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be a sequence of lower semcontinuous proper convex functions such that the sequence $(f_{n}^{*})_{n\in\mathbb{N}}$ satisfies the equicoercivity property
(4.1) $(x_{n}^{*})_{n\in\mathbb{N}}\subset X^{*},\quad\sup_{n\in\mathbb{N}}f_{n}^{*}(x_{n}^{*})<\infty\implies\sup_{n\in\mathbb{N}}\|x_{n}^{*}\|_{\ast}<\infty.$
Then it holds that
$(\Gamma\text{-}\displaystyle\liminf_{n}f_{n})^{*}={\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{*},$
provided that ${\rm dom}\kern 1.19995pt\Gamma\text{-}\displaystyle\limsup_{n}f_{n}\neq\emptyset$.
The following example emphasizes that the condition ${\rm dom}\kern 1.19995pt\Gamma\text{-}\displaystyle\limsup_{n}f_{n}\neq\emptyset$ is necessary for the validity of Theorem 4.1.
Example 4.2.
Consider the sequence of functions $f_{n}(x):=\tfrac{1}{n}|x|+n$. On the one hand, we have
$f(x):=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}(x)=\infty,$
which implies $f^{\ast}\equiv$-$\infty$. On the other hand,
$f_{n}^{\ast}(x^{\ast})=\delta_{[-\tfrac{1}{n},\tfrac{1}{n}]}(x^{\ast})-n,$
and therefore
${\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{\ast}(x^{\ast})=\begin{cases}-\infty&\text{if }x^{\ast}=0,\\[3.00003pt] \infty&\text{if }x^{\ast}\neq 0.\end{cases}$
Hence, $f^{\ast}\neq{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{\ast}$.
In the following example, we show that, in the last corollary, the assumption that $X$ is WCG cannot, in general, be removed for an arbitrary Banach space $X$.
Example 4.3.
Let $X$ be a Banach space, and suppose there exists a sequence of uniformly bounded, $w^{\ast}$-closed, convex sets $C_{n}\subset X^{\ast}$ such that $C_{n}\subset C_{n+1}$ for all $n\in\mathbb{N}$, and
$\operatorname{Ls}_{w^{\ast}}C_{n}\neq C:={\rm cl}\kern 1.19995pt^{w^{\ast}}\!\Big(\bigcup_{n\in\mathbb{N}}C_{n}\Big).$
Define functions $f,f_{n}:X\to\mathbb{R}$, $n\in\mathbb{N}$, by
$f_{n}(x):=\sigma_{C_{n}}(x),\qquad f(x):=\sigma_{C}(x).$
It is straightforward to verify that $\Gamma\text{-}\displaystyle\liminf_{n}f_{n}=f$, which in particular implies
${\rm dom}\kern 1.19995pt(\Gamma\text{-}\displaystyle\limsup_{n}f_{n})\neq\emptyset.$
Moreover, since $f_{n}^{\ast}=\delta_{C_{n}}$, the uniform boundedness of the sets $C_{n}$ immediately yields the equicoercivity property (4.1).
Now, let $x^{\ast}\in C\setminus\operatorname{Ls}_{w^{\ast}}C_{n}$. On the one hand, we have $f^{\ast}(x^{\ast})=0$. On the other hand, suppose that there exists a sequence $x_{n}^{\ast}\to x^{\ast}$ with $\limsup_{n}f^{\ast}_{n}(x^{\ast}_{n})<\infty$. Then the inclusion $x_{n}^{\ast}\in C_{n}$ must hold for all sufficiently large $n\in\mathbb{N}$; otherwise, we would obtain $\limsup_{n}f^{\ast}_{n}(x^{\ast}_{n})=\infty$. However, this would imply $x^{\ast}\in\operatorname{Ls}_{w^{\ast}}C_{n}$, which is a contradiction. Therefore, the only possible alternative is
$\limsup_{n}f_{n}^{\ast}(x_{n}^{\ast})=\infty.$
By the arbitrariness of the sequence $x_{n}^{\ast}\to x^{\ast}$, we conclude that ${\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}(x^{\ast})=\infty$. Consequently, Theorem 4.1 does not hold in this setting.
To illustrate this with a concrete example, let us consider $X=\ell_{\infty}(\mathbb{N})$, and define the sets
$C_{k}:=\left\{(x_{i})_{i\in\mathbb{N}}\in\ell_{1}(\mathbb{N}):x_{j}=0\text{ for all }j\geq k,\ \sum_{j\in\mathbb{N}}|x_{j}|\leq 1\right\}.$
These sets can be regarded as subsets of $X^{\ast}$, using the identification $\ell_{1}(\mathbb{N})\subset X^{\ast}=\ell_{\infty}(\mathbb{N})^{\ast}=\left(\ell_{1}(\mathbb{N})\right)^{\ast\ast}$. It is not difficult to show that each $C_{k}$ is $w^{\ast}$-closed and convex. Furthermore, by Goldstine’s theorem (see, e.g., [43, Theorem 3.96]), we obtain
$C:={\rm cl}\kern 1.19995pt^{w^{\ast}}\left(\bigcup_{n\in\mathbb{N}}C_{n}\right)=\mathbb{B}_{X^{\ast}}:=\{x^{\ast}\in X^{\ast}:\|x^{\ast}\|_{\ast}\leq 1\}.$
Finally, by the Odell–Rosenthal theorem (see, e.g., [59] or [41, Chapter XIII, Theorem 10]), the closed unit ball in $\ell_{1}(\mathbb{N})$ is not weak sequentially dense in the closed unit ball of $X^{\ast}$. Particularly, it implies that
$\operatorname{Ls}_{w^{\ast}}C_{n}\neq C.$
Remark 4.4.
It is worth noting that the equicoercivity property given in (4.1) can be readily verified under the stronger assumption of uniform coercivity, that is, there exist constants $\alpha>0$ and $\beta\in\mathbb{R}$ such that
(4.2) $f_{n}^{\ast}(x^{\ast})\geq\alpha\|x^{\ast}\|+\beta,\quad\text{for all }x^{\ast}\in X^{\ast}\text{ and all }n\in\mathbb{N}.$
By taking the Fenchel conjugate in (4.2) (see, e.g., [28, Fact 4.4.9]), one obtains that this condition is equivalent to the uniform boundedness of the functions in a neighborhood of the origin, which menas, there exist $\rho,\varepsilon>0$ such that
(4.3) $f_{n}(x)\leq\rho,\quad\text{for all }x\in\mathbb{B}_{\varepsilon}(0)\text{ and all }n\in\mathbb{N}.$
Furthermore, by [81, Theorem 2.2.22], this is in turn equivalent to the existence of a convex open neighborhood $U$ of the origin such that, for every $x\in U$, the sequence $(f_{n}(x))_{n\in\mathbb{N}}$ is bounded.
Now, let us recall the following result established first by Mosco in [57] and generalized by Joly in [48]. Nonetheless, a proof can be found in [62, Theorem 2].
Theorem 4.5.
Let $X$ be a Banach space and let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be a sequence of lower semicontinuous proper convex functions. Suppose there exists a bounded sequence $(x_{n}^{\ast})_{n\in\mathbb{N}}$ in $X^{\ast}$ such that the sequence $(f_{n}^{\ast}(x_{n}^{\ast}))_{n\in\mathbb{N}}$ is bounded above. Then the following equality holds:
$({\Gamma(w^{\ast})\text{-}\displaystyle\liminf_{n}}f_{n}^{\ast})^{*}=\Gamma\text{-}\displaystyle\limsup_{n}f_{n}.$
As an immediate consequence of the two preceding results, we obtain a dual formula that relates the epigraphical limit in the strong topology to the epigraphical limits of the corresponding conjugates with respect to the weak topology. The proof of this result is presented in Section 7.
Corollary 4.6.
Let $X$ be a Banach space and let $f,f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be lower semicontinuous proper convex functions. Then,
$f^{\ast}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}f_{n}^{\ast}\Rightarrow f=\Gamma\text{-}\displaystyle\lim_{n}f_{n}.$
Additionally, assume that $X$ is a WCG Banach space and the sequence $(f_{n}^{\ast})_{n\in\mathbb{N}}$ satisfies the equicoercivity property given in (4.1). Then
$f=\Gamma\text{-}\displaystyle\lim_{n}f_{n}\quad\Longleftrightarrow\quad f^{\ast}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}f_{n}^{\ast}.$
Next, we establish a connection between the epigraphical limits of a sequence of convex functions and the convergence of their associated subdifferentials in the sense of (3.2). It is worth noting that the $\|\cdot\|\times w^{\ast}$ sequential convergence of subdifferentials has already been investigated in [80]. More precisely, in [80, Theorem 3.1] the authors proved that if a sequence of proper, lower semicontinuous convex functions Mosco-converges to $f$ and satisfies (4.3), then the corresponding subdifferentials converge with respect to $\|\cdot\|\times w^{\ast}$ in the sense of (3.2).
In what follows, we show that this implication can be strengthened: the uniform boundedness condition in (4.3) is in fact equivalent to the epigraphical convergence of the functions. The proof of this equivalence is deferred to Section 8.
Theorem 4.7.
Let $X$ be a WCG space, and let $f,f_{n}\colon X\colon\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions. Consider the following statements:
  1. (a)
    $f^{\ast}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}f^{\ast}_{n}$.
  2. (b)
    $\partial f_{n}\xrightarrow{G(\|\cdot\|\times w^{\ast})}\partial f$ and there exists $(a,a^{*})\in\partial f$ and a sequence $(a_{n},a_{n}^{*})\in\partial f_{n}$ such that $a_{n}\to a,\;a_{n}^{*}\overset{\ast}{\rightharpoonup}a^{*}\text{ and }f_{n}(a_{n})\to f(a).$
  3. (c)
    $f=\Gamma\text{-}\displaystyle\lim_{n}f_{n}$.
Then, $(a)\Rightarrow(b)\Rightarrow(c)$. Moreover, if the sequence $(f^{\ast}_{n})_{n\in\mathbb{N}}$ satisfies the equicoercivity condition (4.1), then the three statements are equivalent.
Observe that [80, Example 3.2] also serves as a counterexample to Theorem 4.7, showing that $(c)$ does not imply $(b)$ once the WCG hypothesis is dropped.
In the setting of a reflexive Banach space $X$, the roles of $X$ and $X^{\ast}$ can be interchanged in the above result. Hence, the equicoercivity assumption (4.1) can be formulated in the primal space. Moreover, since in this case the $w^{\ast}$-topology on $X^{\ast}$ coincides with the weak topology, we obtain the following corollary as a direct consequence of Theorem 4.7.
Corollary 4.8.
Let $X$ be a reflexive Banach space, and let $f,f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions. Consider the following statements:
  1. (a)
    $f={\Gamma(w)\text{-}\displaystyle\lim_{n}}f_{n}$.
  2. (b)
    $\partial f_{n}\xrightarrow{G(w\times\|\cdot\|_{\ast})}\partial f$ and there exist $(a,a^{*})\in\partial f$ and a sequence $(a_{n},a_{n}^{*})\in\partial f_{n}$ such that $a_{n}\rightharpoonup a$, $a^{\ast}_{n}\to a^{\ast}$ and $f_{n}(a_{n})\to f(a)$.
  3. (c)
    $f^{\ast}=\Gamma\text{-}\displaystyle\lim_{n}f^{\ast}_{n}.$
Then the implication $(a)\Longrightarrow(b)\Longrightarrow(c)$ holds. Moreover, if we assume that the sequence $(f_{n})_{n\in\mathbb{N}}$ satisfies
(4.4) $(x_{n})_{n\in\mathbb{N}}\subset X,\quad\sup_{n\in\mathbb{N}}f_{n}(x_{n})<\infty\implies\sup_{n\in\mathbb{N}}\|x_{n}\|<\infty.$
Then the equivalence $(a)\Longleftrightarrow(b)\Longleftrightarrow(c)$ holds.
Lastly, we get the following result, which relates the $G(\left\lVert\cdot\right\rVert\times w^{\ast})$, the $G(w\times\left\lVert\cdot\right\rVert_{\ast})$ and the $G(\left\lVert\cdot\right\rVert\times\left\lVert\cdot\right\rVert_{\ast})$ convergence of subdifferentials in the reflexive case.
Corollary 4.9.
Let $X$ be a reflexive Banach space, and let $f,f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions. Consider the following statements:
  1. (a)
    $\partial f_{n}\xrightarrow{G(\|\cdot\|\times w^{\ast})}\partial f$ and there exist $(a,a^{*})\in\partial f$ and $(a_{n},a_{n}^{*})\in\partial f_{n}$ such that $a_{n}\to a$, $a_{n}^{*}\overset{\ast}{\rightharpoonup}a^{*}$ and $f_{n}(a_{n})\to f(a)$.
  2. (b)
    $\partial f_{n}\xrightarrow{G(w\times\|\cdot\|_{\ast})}\partial f$ and there exist $(a,a^{*})\in\partial f$ and a sequence $(a_{n},a_{n}^{*})\in\partial f_{n}$ such that $a_{n}\rightharpoonup a$, $a^{\ast}_{n}\to a^{\ast}$ and $f_{n}(a_{n})\to f(a)$.
  3. (c)
    $\partial f_{n}\xrightarrow{G(\left\lVert\cdot\right\rVert\times\left\lVert\cdot\right\rVert_{\ast})}\partial f$ and there exist $(a,a^{*})\in\partial f$ and a sequence $(a_{n},a_{n}^{*})\in\partial f_{n}$ such that $a_{n}\to a$, $a^{\ast}_{n}\to a^{\ast}$ and $f_{n}(a_{n})\to f(a)$.
If the sequence $(f_{n})_{n\in\mathbb{N}}$ satisfies (4.4), then
$(a)\wedge(b)\Longleftrightarrow(c).$
[…]
5. Some Preliminary Results
In this section, we collect several results that will be used in the sequel.
The first lemma states a rule for interchanging upper limits. A proof can be found in [4, Corollary 1.16].
Lemma 5.1.
Let $\{\alpha_{k,n}:k\geq 1,\;n\geq 1\}\subset\mathbb{R}\cup\{-\infty,\infty\}$ be a double indexed sequence of extended real numbers. Then, there is $(n_{k})_{k\in\mathbb{N}}$ with $n_{k}\to\infty$ as $k\to\infty$ such that
$\limsup_{k\to\infty}\alpha_{k,n_{k}}\leq\limsup_{n\to\infty}\limsup_{k\to\infty}\alpha_{k,n}.$
The second lemma provides sufficient conditions for the continuity of the Fenchel conjugate with respect to the norm topology in the dual space. This result follows directly from [28, Corollary 4.4.11].
Lemma 5.2.
Let $f\colon X\to\mathbb{R}\cup\{\infty\}$ be a convex function. If there exist $\alpha>0$, $\beta\in\mathbb{R}$, and $\bar{x}^{\ast}\in X^{\ast}$ such that for all $x\in X$
$f(x)\geq\alpha\|x\|^{2}+\langle\bar{x}^{\ast},x\rangle+\beta,$
then $f^{\ast}$ is continuous with respect to the norm topology.
The next lemma concerns the approximation of a convex proper lower semicontinuous function by its Moreau envelope, together with its continuity properties. We refer to [76, Theorem 3.265] and [28, Theorem 4.4.17] for more details.
Lemma 5.3.
Let $f\colon X\to\mathbb{R}\cup\{\infty\}$ be a proper lower semicontinuous convex function. Then, for every $x\in X$, one has
$f_{\lambda}(x)\nearrow f(x)\quad\text{as }\lambda\to 0,$
and $f_{\lambda}$ is continuous for every $\lambda>0$.
The following lemma establishes an approximation principle for conjugate functions via infimal convolutions. Its proof relies on the same ideas as the classical approximation of functions by infimal convolution, and is therefore omitted.
Lemma 5.4.
Let $f\colon X\to\mathbb{R}\cup\{\infty\}$ be a proper lower semicontinuous convex function, and let $A\subset X$ be a weakly compact, convex, and symmetric set such that the space generated by $A$ is dense in $X$ with the norm topology. In addition, suppose that $f$ is continuous at 0. Then, for every $x^{\ast}\in X^{\ast}$,
$g_{n}(x^{\ast}):=f^{\ast}\,\square\,\sigma_{[0,n]A}(x^{\ast})\nearrow f^{\ast}(x^{\ast})\quad\text{as }n\to\infty,$
and $g_{n}$ is continuous with respect to the Mackey topology for every $n\in\mathbb{N}$.
6. Proof of Theorem 4.1
In this section, we provide a proof of Theorem 4.1. The result is first established for separable Banach spaces and then extended to general WCG Banach spaces by means of separable reduction techniques, combined with the use of projectional skeletons.
We begin by formally establishing a relation that always holds. This is stated in the following lemma.
Lemma 6.1.
Let $X$ be a Banach space, and let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be a sequence of lower semicontinuous proper convex functions. Define $f:=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$. Then, one always has
$f^{\ast}\leq{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{\ast}.$
Proof. Let $x^{\ast}\in X^{\ast}$ and $(x^{\ast}_{n})_{n\in\mathbb{N}}\subset X^{\ast}$ be such that $x^{\ast}_{n}\overset{\ast}{\rightharpoonup}x^{\ast}$. Fix $x\in X$ and a sequence $(x_{n})_{n\in\mathbb{N}}\subset X$ with $x_{n}\to x$. By the definition of the conjugate, for every $n\in\mathbb{N}$ we have
$\langle x^{\ast}_{n},x_{n}\rangle-f_{n}(x_{n})\leq f_{n}^{\ast}(x^{\ast}_{n}).$
Taking the $\limsup$, we obtain
$\limsup_{n\to\infty}\big(\langle x^{\ast}_{n},x_{n}\rangle-f_{n}(x_{n})\big)$ $=\langle x^{\ast},x\rangle+\limsup_{n\to\infty}\big(-f_{n}(x_{n})\big)$
$=\langle x^{\ast},x\rangle-\liminf_{n\to\infty}f_{n}(x_{n})$
$\leq\limsup_{n\to\infty}f_{n}^{\ast}(x^{\ast}_{n}).$
Since the choice of $(x^{\ast}_{n})_{n\in\mathbb{N}}$ with $x^{\ast}_{n}\overset{\ast}{\rightharpoonup}x^{\ast}$ was arbitrary, it follows that
$\langle x^{\ast},x\rangle-\liminf_{n\to\infty}f_{n}(x_{n})\;\leq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{\ast}(x^{\ast}).$
Moreover, since $(x_{n})_{n\in\mathbb{N}}$ was also an arbitrary sequence converging to $x$, we deduce that
$\langle x^{\ast},x\rangle-\Gamma\text{-}\displaystyle\liminf_{n}f_{n}(x)\;\leq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{\ast}(x^{\ast}).$
Equivalently,
$\langle x^{\ast},x\rangle-f(x)\;\leq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{\ast}(x^{\ast}).$
Because this inequality holds for every $x\in X$, we conclude that
$f^{\ast}(x^{\ast})\leq{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{\ast}(x^{\ast}),\text{ for all }x^{\ast}\in X^{\ast},$
which completes the proof of the lemma.
[…]
6.1. Proof of Theorem 4.1 under separability of $X$
In this subsection, we establish Theorem 4.1 in the case where $X$ is separable. Proof. Suppose that $X$ is a separable Banach space and denote $f:=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$. From Lemma 6.1, we already know that
$f^{\ast}\;\leq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{\ast}.$
It remains to prove the reverse inequality in the separable case, namely,
$f^{\ast}\;\geq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{\ast}.$
For that purpose, it is sufficient to show that that for every $x^{*}\in X^{*}$ there exists a sequence $(z_{n}^{*})_{n\in\mathbb{N}}$ such that $z_{n}^{*}\overset{\ast}{\rightharpoonup}x^{*}$ and
(6.1) $f^{*}(x^{*})\geq\limsup_{n}f_{n}^{*}(z_{n}^{*}).$
Let $\{x_{1},x_{2},\dots\}$ be a countable dense subset of $X$. Take $a\in{\rm dom}\kern 1.19995pt(\Gamma\text{-}\displaystyle\limsup_{n}f_{n})$, and let $(a_{n})_{n\in\mathbb{N}}$ be a sequence such that
(6.2) $a_{n}\to a\quad\text{and}\quad\limsup_{n\to\infty}f_{n}(a_{n})=\Gamma\text{-}\displaystyle\limsup_{n}f_{n}(a)<\infty.$
In particular, the sequence $(f_{n}(a_{n}))_{n\in\mathbb{N}}$ may be assumed to be bounded from above.
Let $E_{k}$ be the vector space generated by $\{x_{1},\dots,x_{k}\}$. Define
$A_{k}=\{x\in E_{k}:\|x\|\leq k\},\quad K:=\{a_{n}\}_{n\in\mathbb{N}}\cup\{a\},$
so that $A_{k}\cup K$ is compact with respect to the norm topology. Then set
$B_{k}:=\overline{\operatorname{co}}(A_{k}\cup K).$
Observe that for each $k\in\mathbb{N}$, the set $B_{k}$ is convex and compact (see, e.g., [1, Theorem 5.35]).
We define the following family of functions:
(6.3) $f_{n,k,\lambda}$ $=f_{n}+\delta_{B_{k}}+\frac{\lambda}{2}\|\cdot\|^{2},\quad f_{k,\lambda}=f+\delta_{B_{k}}+\frac{\lambda}{2}\|\cdot\|^{2}.$
Since ${\rm dom}\kern 1.19995ptf_{n}\cap B_{k}\neq\emptyset$ and ${\rm dom}\kern 1.19995ptf\cap B_{k}\neq\emptyset$, all functions defined in (6.3) are proper.
Now, we consider $x^{\ast}\in X^{\ast}$ fixed, but arbitrary. We aim to show (6.1) for some $z^{\ast}_{n}\overset{\ast}{\rightharpoonup}x^{\ast}$. Observe that the case $f^{\ast}(x^{\ast})=\infty$ is trivial. Note also that because of (6.2), $f$ is not identically $\infty$, so the case $f^{\ast}(x^{\ast})=$-$\infty$ is not possible. This said, from now on, we will assume that $f^{\ast}(x^{\ast})\in\mathbb{R}$.
The remainder of the proof is divided into five claims.
Claim 1: For every $k\in\mathbb{N}$ and any $\lambda>0$ we have
(6.4) $\quad f_{k,\lambda}^{*}(x^{*})\geq\limsup_{n}f_{n,k,\lambda}^{*}(x^{*}).$
Indeed, we notice that (6.4) is equivalent to
$\liminf_{n}\inf_{u\in B_{k}}\left\{f_{n}(u)+\frac{\lambda}{2}\|u\|^{2}-\langle x^{*},u\rangle\right\}\geq\inf_{u\in B_{k}}\left\{f(u)+\frac{\lambda}{2}\|u\|^{2}-\langle x^{*},u\rangle\right\}.$
If the left-hand side is equal to $\infty$ the inequality is trivial. Then, let us suppose that the left-hand side is finite. Since $B_{k}$ is compact, we can find $u_{n}\in B_{k}$ such that
$\inf_{u\in B_{k}}\left\{f_{n}(u)+\frac{\lambda}{2}\|u\|^{2}-\langle x^{*},u\rangle\right\}=f_{n}(u_{n})+\frac{\lambda}{2}\|u_{n}\|^{2}-\langle x^{*},u_{n}\rangle$
Furthermore, by definition of limit inferior, we can find a subsequence $u_{n_{j}}\in B_{k}$ such that:
$\lim_{j}\left\{f_{n_{j}}(u_{n_{j}})+\frac{\lambda}{2}\|u_{n_{j}}\|^{2}-\langle x^{*},u_{n_{j}}\rangle\right\}=\liminf_{n}\inf_{u\in B_{k}}\left\{f_{n}(u)+\frac{\lambda}{2}\|u\|^{2}-\langle x^{*},u\rangle\right\}.$
Moreover, again using the compactness of $B_{k}$, by passing to a subsequence, we can assume that $u_{n_{j}}$ converges (strongly) in $B_{k}$ to some $\bar{u}\in B_{k}$. We define the sequence
$u^{\prime}_{k}=\begin{cases}u_{n_{j}}&\text{ if }k=n_{j},\;\text{for some }j\in\mathbb{N}.\\ \bar{u}&\text{ if }k\neq n_{j},\;\forall j\in\mathbb{N}.\end{cases}$
Observe that $u^{\prime}_{k}\to\bar{u}$. Now, using that $f=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$, we obtain that
$\lim_{j}\left\{f_{n_{j}}(u_{n_{j}})+\frac{\lambda}{2}\|u_{n_{j}}\|^{2}-\langle x^{*},u_{n_{j}}\rangle\right\}$ $\geq\liminf_{j}f_{n_{j}}(u_{n_{j}})+\frac{\lambda}{2}\|\bar{u}\|^{2}-\langle x^{*},\bar{u}\rangle$
$\geq\liminf_{j}f_{j}(u^{\prime}_{j})+\frac{\lambda}{2}\|\bar{u}\|^{2}-\langle x^{*},\bar{u}\rangle$
$\geq f(\bar{u})+\frac{\lambda}{2}\|\bar{u}\|^{2}-\langle x^{*},\bar{u}\rangle$
$\geq\inf_{u\in B_{k}}\left\{f(u)+\frac{\lambda}{2}\|u\|^{2}-\langle x^{*},u\rangle\right\}.$
which shows that (6.4) holds.
Claim 2: We have that for all $k\in\mathbb{N}$, all $\lambda>0$, and any $n\in\mathbb{N}$:
(6.5) $f_{n,k,\lambda}^{\ast}=\operatorname{cl}((f_{n}^{\ast})_{\lambda}\square\delta_{B_{k}}^{\ast}),$
where the closure is taken with respect to the weak topology, that is,
[…]
for every $z^{\ast}\in X^{\ast}$
First, by [64, Theorem 3], we have
(6.6) $\left(f_{n}+\frac{\lambda}{2}\left\lVert\cdot\right\rVert^{2}\right)^{\ast}=f^{\ast}_{n}\square\left(\frac{\lambda}{2}\left\lVert\cdot\right\rVert_{\ast}^{2}\right)^{\ast}=\left(f_{n}^{\ast}\right)_{\lambda}.$
where in the last equality we have used that $\left(\frac{\lambda}{2}\left\lVert\cdot\right\rVert^{2}\right)^{\ast}=\frac{1}{2\lambda}\left\lVert\cdot\right\rVert_{\ast}^{2}$. Next, due to the fact that ${\rm dom}\kern 1.19995pt\left(f_{n}+\frac{\lambda}{2}\left\lVert\cdot\right\rVert^{2}\right)\cap{\rm dom}\kern 1.19995pt\delta_{B_{k}}\neq\emptyset$ (recall that $a_{n}\in B_{k}$), we may apply [36, Proposition 4.1.16] to deduce
(6.7) $f_{n,k,\lambda}^{\ast}=\left(f_{n}+\frac{\lambda}{2}\left\lVert\cdot\right\rVert^{2}+\delta_{B_{k}}\right)^{\ast}=\operatorname{cl}\left(\left(f_{n}+\frac{\lambda}{2}\left\lVert\cdot\right\rVert^{2}\right)^{\ast}\square\delta_{B_{k}}^{\ast}\right).$
Hence, combining (6.6) and (6.7), we conclude that (6.5) holds.
Claim 3: There is a positive sequence $(\lambda_{n})_{n\in\mathbb{N}}$ converging to 0, a sequence of natural numbers $(k_{n})_{n\in\mathbb{N}}$ such that $k_{n}\to\infty$ and a sequence $(u^{\ast}_{n})_{n\in\mathbb{N}}\subset X^{\ast}$ with
$\max_{1\leq j\leq n}|\langle x^{\ast}-u^{\ast}_{n},x_{j}\rangle|\leq\frac{1}{n}$
such that
(6.8) $f^{\ast}(x^{\ast})\geq\limsup_{n}(f_{n}^{\ast})_{\lambda_{n}}\square\delta_{B_{k_{n}}}^{\ast}(u^{\ast}_{n}).$
Indeed, since $f_{k,\lambda}\geq f$, it follows that $f^{*}\geq f_{k,\lambda}^{*}$. Combining this with (6.4) and using $\lambda(k):=\frac{1}{k}$, we get
$f^{*}(x^{*})\geq\limsup_{n}f_{n,k,\lambda(k)}^{*}(x^{*}).$
Therefore
$f^{*}(x^{*})\geq\limsup_{k}\limsup_{n}f_{n,k,\lambda(k)}^{*}(x^{*}).$
Now we use the technical result provided in Lemma 5.1 to deduce that there is a sequence of natural numbers $(k_{n})_{n\in\mathbb{N}}$ such that $k_{n}\to\infty$ and
(6.9) $f^{*}(x^{*})\geq\limsup_{n}f_{n,k_{n},\lambda(k_{n})}^{*}(x^{*}).$
Now, using (6.5) and the fact that $f^{\ast}_{n,k,\lambda}$ is proper, we have that
$f_{n,k,\lambda}^{\ast}(x^{\ast})+\frac{1}{n}>{\rm cl}\kern 1.19995pt_{w^{\ast}}((f_{n}^{\ast})_{\lambda}\square\delta_{B_{k}}^{\ast})(x^{\ast}).$
It follows that there is a directed set $I_{n}$ and a net $(u^{\ast}_{i,n,k,\lambda})_{i\in I_{n}}$ convergent to $x^{\ast}$ such that
$f_{n,k,\lambda}^{\ast}(x^{\ast})+\frac{1}{n}>\liminf_{i\in I_{n}}(f_{n}^{\ast})_{\lambda}\square\delta_{B_{k}}^{\ast}(u^{\ast}_{i,n,k,\lambda})$
So we can construct a sequence $(u^{\ast}_{n,k,\lambda})_{n\in\mathbb{N}}\subset X^{\ast}$ such that
$f_{n,k,\lambda}^{\ast}(x^{\ast})+\frac{1}{n}$ $>(f_{n}^{\ast})_{\lambda}\square\delta_{B_{k}}^{\ast}(u^{\ast}_{n,k,\lambda}).$
$\max_{1\leq j\leq n}|\langle x^{\ast}-u^{\ast}_{n,k,\lambda},x_{j}\rangle|$ $\leq\frac{1}{n}.$
We can apply this to $k=k_{n},\;\lambda=\lambda_{n}:=\lambda(k_{n})$, and denoting by $u^{\ast}_{n}:=u_{n,k_{n},\lambda_{n}}^{\ast}$, we combine the latter with (6.9) to obtain that
$f^{\ast}(x^{\ast})\geq\limsup_{n}(f_{n}^{\ast})_{\lambda_{n}}\square\delta_{B_{k_{n}}}^{\ast}(u^{\ast}_{n}),$
with $\max_{1\leq j\leq n}|\langle x^{\ast}-u^{\ast}_{n},x_{j}\rangle|\leq\frac{1}{n}$.
Claim 4: There are bounded sequences $(x_{n}^{\ast})_{n\in\mathbb{N}}$ and $(y^{\ast}_{n})_{n\in\mathbb{N}}$ such that $(f^{\ast}_{n}(y^{\ast}_{n}))_{n\in\mathbb{N}}$ is bounded below, $\left\lVert y^{\ast}_{n}-x^{\ast}_{n}\right\rVert_{\ast}\to 0$ and
(6.10) $f^{\ast}(x^{\ast})\geq\limsup_{n}f_{n}^{\ast}(y^{\ast}_{n})+\frac{1}{2\lambda_{n}}\left\lVert x_{n}^{\ast}-y^{\ast}_{n}\right\rVert_{\ast}^{2}\geq\limsup_{n}f_{n}^{\ast}(y^{\ast}_{n}).$
Indeed, using the inf-convolution of (6.8), we obtain that for each $n\in\mathbb{N}$ there exists some $x^{\ast}_{n}\in X^{\ast}$ such that
$(f_{n}^{\ast})_{\lambda_{n}}\square\delta_{B_{k_{n}}}^{\ast}(u^{\ast}_{n})+\frac{1}{n}\geq(f_{n}^{\ast})_{\lambda_{n}}(x_{n}^{\ast})+\delta_{B_{k_{n}}}^{\ast}(u_{n}^{\ast}-x_{n}^{\ast}).$
Then, using (6.8) and the fact that for every $n\in\mathbb{N}$ there is $y_{n}^{\ast}\in X^{\ast}$ such that
$\frac{1}{n}+\displaystyle\inf_{u^{\ast}\in X^{\ast}}\{f_{n}^{\ast}(u^{\ast})+\frac{1}{2\lambda_{n}}\left\lVert x_{n}^{\ast}-u^{\ast}\right\rVert_{\ast}^{2}\}\geq f_{n}^{\ast}(y^{\ast}_{n})+\frac{1}{2\lambda_{n}}\left\lVert x_{n}^{\ast}-y^{\ast}_{n}\right\rVert_{\ast}^{2},$
we deduce that
(6.11) $f^{\ast}(x^{\ast})$ $\geq\limsup_{n}\left((f_{n}^{\ast})_{\lambda_{n}}(x_{n}^{\ast})+\delta_{B_{k_{n}}}^{\ast}(u_{n}^{\ast}-x_{n}^{\ast})\right)$
$=\limsup_{n}\left(\delta_{B_{k_{n}}}^{\ast}(u_{n}^{\ast}-x_{n}^{\ast})+\inf_{u^{\ast}\in X^{\ast}}\{f_{n}^{\ast}(u^{\ast})+\frac{1}{2\lambda_{n}}\left\lVert x_{n}^{\ast}-u^{\ast}\right\rVert_{\ast}^{2}\}\right)$
$\geq\limsup_{n}\left(\delta_{B_{k_{n}}}^{\ast}(u_{n}^{\ast}-x_{n}^{\ast})+f_{n}^{\ast}(y^{\ast}_{n})+\frac{1}{2\lambda_{n}}\left\lVert x_{n}^{\ast}-y^{\ast}_{n}\right\rVert_{\ast}^{2}\right)$
$\geq\limsup_{n}\left(\delta_{B_{k_{n}}}^{\ast}(u_{n}^{\ast}-x_{n}^{\ast})+f_{n}^{\ast}(y^{\ast}_{n})\right)$
$\geq\limsup_{n}f_{n}^{\ast}(y^{\ast}_{n}).$
Since $f^{\ast}(x^{\ast})$ is finite, $(f_{n}^{\ast}(y^{\ast}_{n}))_{n\in\mathbb{N}}$ is bounded above, and by the property of equicoercivity (4.1), we conclude that $(y_{n}^{\ast})_{n\in\mathbb{N}}$ is bounded.
Furthermore, let us note that $(f_{n}^{\ast}(y^{\ast}_{n}))_{n\in\mathbb{N}}$ is bounded below. In fact, using the sequence $(a_{n})_{n\in\mathbb{N}}$ from (6.2):
$\begin{split}\liminf_{n}f_{n}^{\ast}(y^{\ast}_{n})&\geq\liminf_{n}{\langle y_{n}^{\ast},a_{n}\rangle-f_{n}(a_{n})}\\ &\geq\liminf_{n}{(-\left\lVert y_{n}^{\ast}\right\rVert_{\ast}\left\lVert a_{n}\right\rVert-f_{n}(a_{n}))}\\ &=-\limsup_{n}({\left\lVert y_{n}^{\ast}\right\rVert_{\ast}\left\lVert a_{n}\right\rVert+f_{n}(a_{n})}).\end{split}$
Since the sequences $(f_{n}(a_{n}))_{n\in\mathbb{N}}$ and $(\left\lVert y_{n}^{\ast}\right\rVert\left\lVert a_{n}\right\rVert)_{n\in\mathbb{N}}$ are bounded above, then we get
$\liminf_{n}f_{n}^{\ast}(y^{\ast}_{n})>$-$\infty$
and the sequence $(f_{n}^{\ast}(y^{\ast}_{n}))_{n\in\mathbb{N}}$ is bounded below.
Then, returning to (6.11), since $\lambda_{n}\to 0$ and $(f_{n}^{\ast}(y^{\ast}_{n}))_{n\in\mathbb{N}}$ is bounded below, it follows that $\left\lVert x_{n}^{\ast}-y^{\ast}_{n}\right\rVert_{\ast}\to 0$. Thus, since the sequence $(y_{n}^{\ast})_{n\in\mathbb{N}}$ is bounded, it follows that $(x_{n}^{\ast})_{n\in\mathbb{N}}$ is also bounded.
Claim 5: The sequence $(y^{\ast}_{n})_{n\in\mathbb{N}}$ from (6.10) satisfies $y^{\ast}_{n}\overset{\ast}{\rightharpoonup}x^{\ast}$.
We will show that the sequence $(x^{\ast}_{n})_{n\in\mathbb{N}}$ from (6.10) satisfies $x_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}$. In fact, using (6.11) and that the sequence $(f_{n}^{\ast}(y^{\ast}_{n}))_{n\in\mathbb{N}}$ is bounded below, we obtain that
(6.12) $f^{\ast}(x^{\ast})$ $\geq\limsup_{n}f_{n}^{\ast}(y^{\ast}_{n})+\delta_{B_{k_{n}}}^{\ast}(u_{n}^{\ast}-x_{n}^{\ast})$
$\geq\inf_{n\in\mathbb{N}}f^{\ast}(y^{\ast}_{n})+\limsup_{n}\delta_{A_{k_{n}}}^{\ast}(u_{n}^{\ast}-x_{n}^{\ast}).$
Note that by [36, Proposition 4.1.20], we have
$\begin{split}\delta_{A_{k_{n}}}^{\ast}&=(\delta_{E_{k_{n}}}+\delta_{\{x\in X:\left\lVert x\right\rVert\leq k_{n}\}})^{\ast}=\delta_{E_{k_{n}}}^{\ast}\square\delta_{\{x\in X:\left\lVert x\right\rVert\leq k_{n}\}}^{\ast}=\delta_{E_{k_{n}}^{\perp}}\square k_{n}\left\lVert\cdot\right\rVert_{\ast}=k_{n}d(x^{\ast},E_{k_{n}}^{\perp}),\end{split}$
where $d(x^{\ast},E_{k_{n}}^{\perp})$ denotes the distance from $x^{\ast}$ to $E_{k_{n}}^{\perp}$.
Using this and (6.12), we obtain that
$f^{\ast}(x^{\ast})-\inf_{n\in\mathbb{N}}f^{\ast}(y^{\ast}_{n})\geq\displaystyle\limsup_{n}{k_{n}d(u_{n}^{\ast}-x_{n}^{\ast},E_{k_{n}}^{\perp})}.$
Since $k_{n}\to\infty$, it follows that
$d(u_{n}^{\ast}-x_{n}^{\ast},E_{k_{n}}^{\perp})\to 0.$
Then, given $\varepsilon>0$, there is $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ there is $h^{\ast}_{n}\in E_{k_{n}}^{\perp}$ such that
(6.13) $\left\lVert u^{\ast}_{n}-x^{\ast}_{n}-h^{\ast}_{n}\right\rVert_{\ast}\leq\varepsilon.$
Let $\bigcup_{k\in\mathbb{N}}E_{k}, DD$ is dense. Let $x\in D$; Since $D$ is an increasing union, it follows that $x\in E_{k_{n}}$ for sufficiently large $n$.
We have that $\langle u_{n}^{\ast}-x_{n}^{\ast},x\rangle=\langle u_{n}^{\ast}-x_{n}^{\ast}-h^{\ast}_{n},x\rangle$.
Consequently, $\lvert\langle u_{n}^{\ast}-x_{n}^{\ast},x\rangle\rvert\leq\left\lVert u^{\ast}_{n}-x^{\ast}_{n}-h^{\ast}_{n}\right\rVert_{\ast}\left\lVert x\right\rVert\leq\varepsilon\left\lVert x\right\rVert$. Which, by (6.13), implies that
$\langle u_{n}^{\ast}-x_{n}^{\ast},x\rangle\to 0\text{ for all }x\in D.$
Let us prove that $x_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}$. Since $(x_{n}^{\ast})_{n\in\mathbb{N}}$ is norm bounded and $\{x_{j}:j\in\mathbb{N}\}$ is dense in $X$, all we need to proof is that
$\langle x^{\ast}_{n}-x^{\ast},x_{j}\rangle\to 0\text{ for all }j\in\mathbb{N}.$
So let $j\in\mathbb{N}$ and $j\geq n$, then
$|\langle x^{\ast}_{n}-x^{\ast},x_{j}\rangle|=|\langle x^{\ast}_{n}-u^{\ast}_{n},x_{j}\rangle+\langle u^{\ast}_{n}-x^{\ast},x_{j}\rangle|\leq|\langle x^{\ast}_{n}-u^{\ast}_{n},x_{j}\rangle|+\frac{1}{n}.$
It follows that $|\langle x^{\ast}_{n}-x^{\ast},x_{j}\rangle|\to 0$ and this is valid for every $j\in\mathbb{N}$, hence $x_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}$.
Finally, since $\left\lVert x_{n}^{\ast}-y^{\ast}_{n}\right\rVert_{\ast}\to 0$, it is concluded that
(6.14) $y_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}.$
[…]
6.2. Separable reduction and rich families
We say that a property $\mathcal{P}$ conecerning a Banach space $X$, is separably reducible (or separably determined) if there exists a family $\mathcal{F}$ of separable subspaces of $X$ such that $\cup\left\{F:F\in\mathcal{F}\right\}=X$ and the property $\mathcal{P}$ holds in $X$ if and only if for every $F\in\mathcal{F}$ the property $\mathcal{P}$ holds at $F$.
Separable reducibility is a common tool in functional analysis, as it often leads to substantial simplifications and, in some cases, provides the only viable approach to proving certain results. A simple example of a separably reducible property is continuity of a mapping. For a list of some less trivial properties that are separable determined in analysis, we refer the reader to [44]. There are special families of separable subspaces of a given Banach space that allow to implement a separable reduction, that is, the possibility to “lift” a property from separable subspaces to the entire space, since they allow dealing with countable many different properties. These special families are called rich families, and they were introduced by J.M Borwein and W.B Moors in [26]. Rich families were deeply discussed in connection with a number of properties related to differentiability and subdifferentiability in [53] and [45].
Definition 6.2.
Let $Z$ be a (rather) non-separable metric (or normed) space. Denote by $\mathcal{S}(Z)$ the family of all closed separable subsets (or closed separable subspaces) of $Z$. A subfamily $R\subset\mathcal{S}(Z)$ is called rich if:
  1. (i)
    For every $S\in\mathcal{S}(Z)$, there exists $Y\in R$ such that $S\subset Y$.
  2. (ii)
    Whenever $Y_{1}\subset Y_{2}\subset\cdots\in R$, then $Y:=\overline{\displaystyle\bigcup_{n\in\mathbb{N}}Y_{n}}\in R$.
The main property and the power is showed in the following fundamental fact (see [26, Proposition 1.1]).
Proposition 6.3.
The intersection of countably many rich families of a given partial order is (not only non-empty but also) rich.
Let $k\in\mathbb{N}$ be greater than $1$, and let $X_{1},\dots,X_{k}$ be Banach spaces. A block is any product of the form $Y_{1}\times\cdots\times Y_{k}$. A block-family in $\mathcal{S}(X_{1}\times\cdots\times X_{k})$ (or simply in $X_{1}\times\cdots\times X_{k}$) is any family consisting of such blocks, where $Y_{i}\in\mathcal{S}(X_{i})$ for $i=1,\dots,k$.
In the special case $k=2$, we refer to blocks as rectangles and to block-families as rectangle-families. We denote by $\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X_{1}\times X_{2})$ the maximal rectangle-family in $\mathcal{S}(X_{1}\times X_{2})$. Clearly, $\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X_{1}\times X_{2})$ is a rich family in $\mathcal{S}(X_{1}\times X_{2})$.
6.3. Projectional skeletons
Now, let us introduce the notion of projectional skeletons, proposed in [50]. This tool has proved to be both efficient and prominent in characterizing properties of nonseparable spaces by means of seemingly simpler techniques that rely on the separability of the space itself, or on the separability of certain associated function spaces.
Definition 6.4.
A projectional skeleton in the (rather non-separable) Banach space $(X,\|\cdot\|)$ is a family of linear bounded projections $(P_{s}:s\in\Gamma)$ on $X$, indexed by a directed and $\sigma$-complete set $(\Gamma,\leq)$ such that
  1. (i)
    $P_{s}(X)$ is separable for every $s\in\Gamma$,
  2. (ii)
    $X=\displaystyle\bigcup_{s\in\Gamma}P_{s}(X)$,
  3. (iii)
    $P_{t}\circ P_{s}=P_{s}=P_{s}\circ P_{t}$ whenever $s,t\in\Gamma$ and $s\leq t$, and
  4. (iv)
    Given a sequence $s_{1}\leq s_{2}\leq\cdots$ in $\Gamma$, we have
    $P_{s}(X)=\overline{\bigcup_{n\in\mathbb{N}}P_{s_{n}}(X)},$
    where $s:=\displaystyle\sup_{n\in\mathbb{N}}s_{n}$.
For $r\geq 1$, we say that $(P_{s}:s\in\Gamma)$ is an $r$-projectional skeleton if it is a projectional skeleton and
$\|P_{s}(x)\|\leq r\|x\|\text{ for all }s\in\Gamma\text{ and }x\in X.$
In addition, we say that a projectional skeleton $(P_{s}:s\in\Gamma)$ is commutative if
$P_{s}\circ P_{t}=P_{t}\circ P_{s},\text{ for all }s,t\in\Gamma.$
Let $A\subset X$ be a nonempty bounded set. We consider the pseudometric $\rho_{A}$ on $X^{*}$ by
$\rho_{A}(x_{1}^{*},x_{2}^{*}):=\sup\left\{\big|\langle x_{1}^{*}-x_{2}^{*},a\rangle\big|:a\in A\right\},\qquad x_{1}^{*},x_{2}^{*}\in X^{*}.$
Given $\varepsilon\geq 0$, a projectional skeleton $(P_{s}:s\in\Gamma)$ in $X$ is called AA-$\varepsilon$-shrinking if, for every increasing sequence $\gamma_{1}\leq\gamma_{2}\leq\cdots$ in $\Gamma$ and for every $x^{*}\in X^{*}$, we have
$\limsup_{j}\rho_{A}\big(P_{\gamma_{j}}^{*}(x^{*}),P_{\sup\gamma_{i}}^{*}(x^{*})\big)\leq\varepsilon\|x^{*}\|.$
If $\varepsilon=0$, we simply say that the skeleton is $A$-shrinking.
The following is a characterization of weakly compactly generated Banach spaces with projectional skeletons, as presented in [46].
Theorem 6.5.
For a Banach space $(X,\|\cdot\|)$, the following are equivalent:
  • (i)
    $X$ is weakly compactly generated.
  • (ii)
    There exist a bounded closed symmetric convex and linearly dense set $A\subset X$, and a (commutative 1-) projectional skeleton $(P_{\gamma}:\gamma\in\Gamma)$ on $(X,\|\cdot\|)$ (with $\bigcup_{\gamma\in\Gamma}P_{\gamma}^{*}(X^{*})=X^{*}$), which is $A$-shrinking and satisfies that $P_{\gamma}(A)\subset A$ for every $\gamma\in\Gamma$.
Moreover, the index set of the skeleton $\Gamma$ may be assumed to be a rich family in $\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{*})$ such that for every $\gamma:=V\times Y\in\Gamma$, we have that $X=V\oplus Y^{\perp}$ and $P_{\gamma}$ is the projection of $X$ onto $V$, i.e, with range $V$ and kernel $Y^{\perp}$, and also $P^{\ast}_{\gamma}(X^{\ast})=\overline{Y}^{w^{\ast}}$.
[…]
6.4. Preliminary Results on Separable Reduction Techniques
To address the case where $X$ is nonseparable, we first establish the existence of suitable rich families. These tools will enable us to extend Theorem 4.1, previously proved under the separability assumption, to the general setting of WCG Banach spaces.
It is important to note that all the rich families constructed in this section will be rectangle rich families, that is, subsets of $\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ even when certain properties concern only the primal or the dual space individually. This convention simplifies the analysis, as it allows us to work with families that simultaneously satisfy conditions in both the primal and dual spaces by considering their intersections.
The next two lemmas are formulated in the framework of general Banach spaces. They show that certain properties of functions admit separable reduction.
The first result concerns the nonemptiness of the domain of the upper epilimt, in order to show that the hyphoteses ${\rm dom}\kern 1.19995pt(\Gamma\text{-}\displaystyle\limsup_{n}f_{n}))\neq\emptyset$ of Theorem 4.1 is true if we restrinct each $f_{n}$ to certain separable subspaces.
Proposition 6.6.
Let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be a sequence of functions such that ${\rm dom}\kern 1.19995pt\Gamma\text{-}\displaystyle\limsup_{n}f_{n}\neq\emptyset$, then there is a rich family $\mathcal{R}\subset\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ such that for every $V\times Y\in\mathcal{R}$, the sequence $(f_{n}|_{V})_{n\in\mathbb{N}}$ satisfies ${\rm dom}\kern 1.19995pt(\Gamma\text{-}\displaystyle\limsup_{n}(f_{n}|_{V}))\neq\emptyset$.
[…]
The second result establishes a separable reduction principle: the values of the conjugate of a function can be recovered from the conjugates of its restrictions to suitable separable subspaces.
Proposition 6.7.
Let $X$ be a Banach space and let $f\colon X\to\mathbb{R}\cup\{\infty\}$ be a proper lower semicontinuous convex function. Then there exists a rectangle rich family $\mathcal{R}\subset\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ such that, for every $V\times Y\in\mathcal{R}$, one has
(6.15) $(f|_{V})^{\ast}(x^{\ast}|_{V})=f^{\ast}(x^{\ast}),\qquad\text{for all }x^{\ast}\in Y.$
Proof. First, observe that by the definition of the conjugate the following inequality holds whenever $V\subset W$ and $x^{\ast}\in X^{\ast}$:
$(f|_{V})^{\ast}(x^{\ast}|_{V})\leq(f|_{W})^{\ast}(x^{\ast}|_{W}).$
Now, assume that there exist $\alpha>0$, $\beta\in\mathbb{R}$, and $\bar{x}^{\ast}\in X^{\ast}$ such that
(6.16) $f(x)\geq\alpha\|x\|^{2}+\langle\bar{x}^{\ast},x\rangle+\beta,\qquad\forall x\in X.$
From Lemma 5.2, this implies that $(f|_{V})^{\ast}$ is continuous for every $V\subset X$ such that $V\cap{\rm dom}\kern 1.19995ptf\neq\emptyset$.
Consider
$\mathcal{R}:=\bigl\{V\times Y\in\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast}):(f|_{V})^{\ast}(x^{\ast}|_{V})=f^{\ast}(x^{\ast})\;\;\forall x^{\ast}\in Y\bigr\}.$
We shall prove that $\mathcal{R}$ is cofinal. Fix any $Z\in\mathcal{S}(X\times X^{\ast})$. Since $\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ is rich, there exists $V_{0}\times Y_{0}\in\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ such that $Z\subset V_{0}\times Y_{0}$. Let $C_{0}$ be a countable dense subset of $Y_{0}$, and let $S_{0}\subset X$ be a countable set such that, for every $x^{\ast}\in C_{0}$,
$f^{\ast}(x^{\ast})=\sup_{x\in S_{0}}\bigl\{\langle x^{\ast},x\rangle-f(x)\bigr\}.$
Define
$V:=\overline{\operatorname{sp}}(V_{0}\cup S_{0}),\qquad Y:=Y_{0}.$
Clearly, $V\times Y\in\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$. We now show that $V\times Y\in\mathcal{R}$. Fix $x^{\ast}\in Y$, and let $(x^{\ast}_{n})\subset C_{0}$ be a sequence converging to $x^{\ast}$. By construction, for all $n\in\mathbb{N}$ we have
$(f|_{V})^{\ast}(x^{\ast}_{n}|_{V})=f^{\ast}(x^{\ast}_{n}).$
Using the continuity of $f^{\ast}$ and $(f|_{V})^{\ast}$, we conclude that
$(f|_{V})^{\ast}(x^{\ast}|_{V})=f^{\ast}(x^{\ast}).$
Thus $\mathcal{R}$ is nonempty and cofinal.
Next, we prove that $\mathcal{R}$ is $\sigma$-complete. Let $(V_{n}\times Y_{n})_{n\in\mathbb{N}}$ be an increasing sequence in $\mathcal{R}$. We must verify that
$\overline{\bigcup_{n\in\mathbb{N}}V_{n}\times Y_{n}}\in\mathcal{R}.$
Since $(V_{n}\times Y_{n})_{n\in\mathbb{N}}$ is increasing, it follows that
$\overline{\bigcup_{n\in\mathbb{N}}V_{n}\times Y_{n}}=\overline{\bigcup_{n\in\mathbb{N}}V_{n}}\times\overline{\bigcup_{n\in\mathbb{N}}Y_{n}}.$
Denote $V:=\overline{\bigcup_{n\in\mathbb{N}}V_{n}}$ and $Y:=\overline{\bigcup_{n\in\mathbb{N}}Y_{n}}$. As $\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ is rich, we have $V\times Y\in\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$. Now, let $x^{\ast}\in Y$ and choose a sequence $(x^{\ast}_{n})_{n\in\mathbb{N}}\subset\bigcup_{n\in\mathbb{N}}Y_{n}$ converging to $x^{\ast}$. Then there exists a subsequence $(Y_{n_{k}})_{k\in\mathbb{N}}$ such that
$x^{\ast}_{k}\in Y_{n_{k}}\text{ and }(f|_{V})^{\ast}(x^{\ast}_{k}|_{V})=(f|_{V_{n_{k}}})^{\ast}(x^{\ast}_{k}|_{V_{n_{k}}})=f^{\ast}(x^{\ast}_{k}),\qquad\text{ for all }k\in\mathbb{N}.$
By continuity of the involved functions, we obtain
$(f|_{V})^{\ast}(x^{\ast}|_{V})=f^{\ast}(x^{\ast}),$
which shows that $V\times Y\in\mathcal{R}$. Therefore, $\mathcal{R}$ is rich.
In the general case where (6.16) does not necessarily hold, define for each $\lambda>0$ the function
$f^{\lambda}(x):=f(x)+\tfrac{\lambda}{2}\|x\|^{2},$
which satisfies (6.16). It follows that, for every $\lambda>0$, the family
$\mathcal{R}_{\lambda}:=\bigl\{V\times Y\in\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast}):(f^{\lambda}|_{V})^{\ast}(x^{\ast}|_{V})=(f^{\lambda})^{\ast}(x^{\ast})\;\;\forall x^{\ast}\in Y\bigr\}$
is rich. Using the known identity $(f^{\lambda}|_{V})^{\ast}=((f|_{V})^{\ast})_{\lambda}$ (the Moreau envelope in $V^{\ast}$, we obtain
$\mathcal{R}_{\lambda}=\bigl\{V\times Y\in\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast}):((f|_{V})^{\ast})_{\lambda}(x^{\ast}|_{V})=(f^{\ast})_{\lambda}(x^{\ast})\;\;\forall x^{\ast}\in Y\bigr\}.$
Finally, set
$\mathcal{R}:=\bigcap_{n\in\mathbb{N}}\mathcal{R}_{\frac{1}{n}}.$
This family is rich. Since $((f|_{V})^{\ast})_{\lambda}\nearrow(f|_{V})^{\ast}$ pointwise as $\lambda\to 0$ (see Lemma 5.3), it follows that $\mathcal{R}$ is the desired family.
[…]
The next proposition establishes that, when the space $X$ is weakly compactly generated, the equality given in (6.15) extends to the weak-closure of the space $Y$. Formally, we state the following result.
Proposition 6.8.
Let $X$ be a WCG Banach space and let $f\colon X\to\mathbb{R}\cup\{\infty\}$ be a proper lower semicontinuous convex function. Then there exists a rectangle rich family $\mathcal{R}\subset\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ such that, for every $V\times Y\in\mathcal{R}$, one has
(6.17) $(f|_{V})^{\ast}(x^{\ast}|_{V})=f^{\ast}(x^{\ast}),\qquad\text{for all }x^{\ast}\in\overline{Y}^{w^{\ast}}.$
Proof. First, assume that $f$ is continuous at the origin.
Since $X$ is WCG, we may consider a weakly compact, convex, and symmetric set $A$ that is linearly dense in the norm topology. It then follows that
$\overline{\bigcup_{n\in\mathbb{N}}[0,n]A}=X.$
Because $f$ is continuous at the origin, it is also continuous at a point of ${\rm dom}\kern 1.19995pt(\delta_{[0,n]A})$. Define
$g_{n}:=f+\delta_{[0,n]A},$
which is a sequence of proper lower semicontinuous convex functions. Let $\mathcal{R}_{n}$ be the rich family associated with $g_{n}$ as in Proposition 6.7, and define the rich family
$\mathcal{R}^{\prime}:=\bigcap_{n\in\mathbb{N}}\mathcal{R}_{n}.$
We will prove that for every $V\times Y\in\mathcal{R}^{\prime}$, the equality (6.17) holds.
Indeed, let $V\times Y\in\mathcal{R}$, and let $x^{\ast}\in\overline{Y}^{w^{\ast}}$. By the Mackey–Arens theorem (see, e.g., [46, Lemma 19]), there exists a net $(x^{\ast}_{\alpha})\subset Y$ converging to $x^{\ast}$ with respect to the Mackey topology, that is, $x^{\ast}_{\alpha}\overset{\mu^{\ast}}{\to}x^{\ast}$. Moreover, from [36, Proposition 4.1.20], we have
$g_{n}^{\ast}=f^{\ast}\,\square\,\sigma_{[0,n]A}.$
Observe that
$g_{n}|_{V}=f|_{V}+(\delta_{[0,n]A})|_{V}=f|_{V}+\delta_{V\cap[0,n]A},$
so that
$(g_{n}|_{V})^{\ast}=(f|_{V})^{\ast}\,\square\,\sigma_{V\cap[0,n]A}.$
Furthermore, it is easy to prove that for all $y^{\ast},z^{\ast}\in X^{\ast}$, and all $n\in\mathbb{N}$ we have
(6.18) $|g^{\ast}_{n}(y^{\ast})-g^{\ast}_{n}(z^{\ast})|\leq\sigma_{[0,n]A}(y^{\ast}-z^{\ast}).$
Similarly,
(6.19) $|(g_{n}|_{V})^{\ast}(y^{\ast}|_{V})-(g_{n}|_{V})^{\ast}(z^{\ast}|_{V})|\leq\sigma_{V\cap[0,n]A}(y^{\ast}|_{V}-z^{\ast}|_{V})\leq\sigma_{[0,n]A}(y^{\ast}-z^{\ast}).$
Since $x^{\ast}_{\alpha}\overset{\mu^{\ast}}{\to}x^{\ast}$ and $[0,n]A$ is weakly compact, convex, and symmetric, it follows from (6.18) and (6.19) that
$(g_{n}|_{V})^{\ast}(x_{\alpha}^{\ast})\to(g_{n}|_{V})^{\ast}(x^{\ast})\quad\text{and}\quad g^{\ast}_{n}(x_{\alpha}^{\ast})\to g^{\ast}_{n}(x^{\ast}).$
On the other hand, by the property of the rich family (see Proposition 6.7),
$(g_{n}|_{V})^{\ast}(x_{\alpha}^{\ast})=g^{\ast}_{n}(x_{\alpha}^{\ast}).$
Hence, we conclude that
(6.20) $(g_{n}|_{V})^{\ast}(x^{\ast})=g^{\ast}_{n}(x^{\ast}),\text{ for all }n\in\mathbb{N}.$
Since $f$ is continuous at the origin, $f|_{V}$ is also continuous. Taking the limit in (6.20) as $n\to\infty$, and applying Lemma 5.4, we obtain (6.17).
Now, for the general case, let $\hat{f}$ denote the Moreau envelope of $f$, i.e., $\hat{f}:=f_{\lambda}$ for some $\lambda>0$. Then $\hat{f}$ is continuous. Let $\mathcal{R}_{\lambda}$ be the rich family associated with $\hat{f}$ as constructed above, and let $\mathcal{A}$ be the rich family associated with the function $h(x):=\tfrac{1}{2}\|x\|^{2}$. Define
$\mathcal{R}:=\mathcal{R}_{\lambda}\cap\mathcal{A}.$
For every $V\times Y\in\mathcal{R}$, we have
$(f|_{V})^{\ast}(x^{\ast}|_{V})+\frac{\lambda}{2}\|x^{\ast}|_{V}\|_{V^{\ast}}^{2}=f^{\ast}(x^{\ast})+\frac{\lambda}{2}\|x^{\ast}\|^{2},\quad\forall x^{\ast}\in\overline{Y}^{w^{\ast}},$
where $\|x^{\ast}|_{V}\|_{V^{\ast}}$ denotes the dual norm on $V^{\ast}$ induced by $\|\cdot\|$ on $V$. By the properties of $\mathcal{A}$, it follows that
$\|x^{\ast}|_{V}\|_{V^{\ast}}=\|x^{\ast}\|_{\ast}.$
Thus, (6.17) holds.
[…]
Despite the fact that the previous result will be primarily used in the proof of Theorem 4.1, it is worth noting that it also leads to interesting insights into the structure of WCG Banach spaces. In particular, it provides a simple isometry between separable subspaces and the weak-closures of separable subspaces in the dual. This result is in the same spirit as [37, Theorem 2.3], but formulated for WCG spaces rather than Asplund spaces. Formally, we obtain the following result.
Corollary 6.9.
Let $X$ be a WCG Banach space. Then there exists a rectangle rich family $\mathcal{R}\subset\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ such that for every $V\times Y\in\mathcal{R}$ and every $x^{\ast}\in V^{\ast}$, there exists $y^{\ast}\in\overline{Y}^{w^{\ast}}$ satisfying
$y^{\ast}|_{V}=x^{\ast}\text{ and }\|y^{\ast}\|_{X^{\ast}}=\|x^{\ast}\|_{V^{\ast}}.$
In other words, for every $V\times Y\in\mathcal{R}$: $\overline{Y}^{w^{\ast}}\ni y^{\ast}\to y^{\ast}|_{V}\in V^{\ast}$ is a surjective isometry.
Proof. Let $\mathcal{R}_{1}$ be the rich family obtained by applying Proposition 6.8 to the function $\delta_{\mathbb{B}}$, where $\mathbb{B}:=\{x\in X:\left\lVert x\right\rVert\leq 1\}$. Let $\mathcal{R}_{2}$ be the rich family from Theorem 6.5, and recall that for $\gamma:=V\times Y\in\mathcal{R}_{2}$ we have $P^{\ast}_{\gamma}(X^{\ast})=\overline{Y}^{w^{\ast}}$. Now set $\mathcal{R}:=\mathcal{R}_{1}\cap\mathcal{R}_{2}$ and consider $\gamma:=V\times Y\in\mathcal{R}$.
Furthermore, let us introduce the following notation to view the projection $P_{\gamma}$ as a mapping from $X$ to $V$:
(6.21) $\tilde{P}_{\gamma}\colon X\to V,\qquad\tilde{P}_{\gamma}(x):=P_{\gamma}(x).$
In this terminology, the adjoint operator of $\tilde{P}_{\gamma}$ is the mapping $\tilde{P}_{\gamma}^{\ast}\colon V^{\ast}\to X^{\ast}$. With this notation, we obtain
$\tilde{P}^{\ast}_{\gamma}(V^{\ast})=P^{\ast}_{\gamma}(X^{\ast}).$
Therefore, for every $x^{\ast}\in V^{\ast}$ there exists $y^{\ast}:=\tilde{P}_{\gamma}^{\ast}(x^{\ast})$ such that $y^{\ast}|_{V}=x^{\ast}$. This shows that the mapping
$\overline{Y}^{w^{\ast}}\ni y^{\ast}\;\longmapsto\;y^{\ast}|_{V}\in V^{\ast}$
is surjective.
Finally, since $\mathcal{R}\subset\mathcal{R}_{1}$, we have for every $y^{\ast}\in\overline{Y}^{w^{\ast}}$ that
$\left\lVert y^{\ast}\right\rVert_{X^{\ast}}=\left\lVert y^{\ast}|_{V}\right\rVert_{V^{\ast}}.$
Hence, the mapping is an isometry, which completes the proof.
[…]
The next proposition shows that the equicoercivity property (4.1) is separably reducible.
Proposition 6.10.
Let $X$ be a WCG Banach space. Let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be a sequence of proper lower semicontinuous convex function satisfying the equicoercivity property (4.1). Then there exists a rectangle rich family $\mathcal{R}\subset\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ such that for every $V\times Y\in\mathcal{R}$ and every sequence $(v^{\ast}_{n})_{n\in\mathbb{N}}\subset V^{\ast}$
[…]
Proof. For each $n\in\mathbb{N}$, let $\mathcal{R}_{n}$ be the rich family from Proposition 6.8 associated with $f_{n}$, and let $\mathcal{F}$ be the rich family from Corollary 6.9. Define
$\mathcal{R}:=\mathcal{F}\cap\bigcap_{n\in\mathbb{N}}\mathcal{R}_{n},$
which is itself a rich family by Proposition 6.3. Let $\gamma:=V\times Y\in\mathcal{R}$ and $(v^{\ast}_{n})\subset V^{\ast}$ with
$\sup_{n\in\mathbb{N}}(f_{n}|_{V})^{\ast}(v^{\ast}_{n})<\infty.$
On the one hand, by the property of $\mathcal{F}$, there exists $y^{\ast}_{n}\in Y$ such that $y^{\ast}_{n}|_{V}=v^{\ast}_{n}$ and $\|v^{\ast}_{n}\|=\|y^{\ast}_{n}\|$. On the other hand, by the properties of each $\mathcal{R}_{n}$, we have
$f_{n}^{\ast}(y^{\ast}_{n})=(f_{n}|_{V})^{\ast}(y^{\ast}_{n}|_{V})=(f_{n}|_{V})^{\ast}(v^{\ast}_{n}).$
Therefore,
$\sup_{n\in\mathbb{N}}f_{n}^{\ast}(y^{\ast}_{n})<\infty.$
[…]
The final result of this section provides an intermediate step between Theorem 4.1 and its generalization. In particular, it shows that there is a sufficiently large family of separable subspaces where the conclusion of Theorem 4.1 holds.
Proposition 6.11.
let $X$ be a WCG Banach space, and let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$ be a sequence of proper lower semicontinuous convex functions such that the sequence $f^{\ast}_{n}$ satisfies (4.1) and such that ${\rm dom}\kern 1.19995pt(\Gamma\text{-}\displaystyle\limsup_{n}f_{n})\neq\emptyset$. Then there is a rectangle rich family $\mathcal{R}\subset\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ such that for every $V\times Y\in\mathcal{R}$ the following holds
$(\Gamma\text{-}\displaystyle\liminf_{n}(f_{n}|_{V}))^{*}(v^{\ast})={\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}(f_{n}|_{V})^{*}(v^{\ast}),\text{ for all }v^{\ast}\in V^{\ast}.$
Proof. Let $\mathcal{A}_{1}$ be the rich family obtained in Proposition 6.10, and let $\mathcal{A}_{2}$ be the rich family from Proposition 6.6. Set
$\mathcal{R}:=\mathcal{A}_{1}\cap\mathcal{A}_{2}.$
[…]
6.5. Proof of Theorem 4.1 in the general case
We now have all the necessary ingredients to extend Theorem 4.1 from the separable case to the general setting of WCG Banach spaces.
Proof. Let $\mathcal{R}\subset\mathcal{S}_{{}_{\sqsubset\!\!\!\sqsupset}}(X\times X^{\ast})$ be a rectangle rich family that simultaneously satisfies Theorem 6.5, Proposition 6.8, Corollary 6.9, Proposition 6.10, and Proposition 6.11.
Denote $f:=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$ and fix $x^{\ast}\in X^{\ast}$. Let $\gamma:=V\times Y\in\mathcal{R}$ be such that $x^{\ast}\in Y$. By Lemma 6.1, we may restrict to the case $f^{\ast}(x^{\ast})<\infty$.
We first observe that
$f|_{V}\;\leq\;\Gamma\text{-}\displaystyle\liminf_{n}f_{n}|_{V}.$
Consequently,
$f^{\ast}(x^{\ast})\;\geq\;(f|_{V})^{\ast}(x^{\ast}|_{V})\;\geq\;\bigl(\Gamma\text{-}\displaystyle\liminf_{n}(f_{n}|_{V})\bigr)^{*}(x^{\ast}|_{V})\;=\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}(f_{n}|_{V})^{*}(x^{\ast}|_{V}),$
where the last equality follows from the properties of the rich family $\mathcal{R}$.
Now fix $\varepsilon>0$. By the definition of ${\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}(f_{n}|_{V})^{*}$, there exists a sequence $(z_{n}^{\ast})\subset V^{\ast}$ with $z_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}|_{V}$ such that
${\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}(f_{n}|_{V})^{*}(x^{\ast}|_{V})>\limsup_{n}(f_{n}|_{V})^{*}(z_{n}^{\ast})-\varepsilon.$
By Proposition 6.8 and Corollary 6.9, there exist $x_{n}^{\ast}\in\overline{Y}^{w^{\ast}}$ such that $x_{n}^{\ast}|_{V}=z_{n}^{\ast}$ and
$(f_{n}|_{V})^{*}(z_{n}^{\ast})=f_{n}^{*}(x_{n}^{\ast}).$
Moreover, $\tilde{P}^{\ast}_{\gamma}(z_{n}^{\ast})=x_{n}^{\ast}$ and $\tilde{P}^{\ast}_{\gamma}(x^{\ast}|_{V})=x^{\ast}$, where $\tilde{P}_{\gamma}$ is defined in (6.21).
We now show that $x_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}$. Indeed, for any $x\in X$,
$\bigl|\langle x_{n}^{\ast}-x^{\ast},x\rangle\bigr|$ $=\bigl|\langle\tilde{P}^{\ast}_{\gamma}(z_{n}^{\ast})-\tilde{P}^{\ast}_{\gamma}(x^{\ast}|_{V}),x\rangle\bigr|$
$=\bigl|\langle z_{n}^{\ast}-x^{\ast}|_{V},\tilde{P}_{\gamma}(x)\rangle\bigr|.$
Since $\tilde{P}_{\gamma}(x)=P_{\gamma}(x)\in V$ and $z_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}|_{V}$ in $V^{\ast}$, it follows that $\langle x_{n}^{\ast}-x^{\ast},x\rangle\to 0$.
Hence $x_{n}^{\ast}\overset{\ast}{\rightharpoonup}x^{\ast}$. Therefore,
$\limsup_{n}f_{n}^{*}(x_{n}^{\ast})\;\geq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{*}(x^{\ast}).$
Consequently,
$f^{\ast}(x^{\ast})\;\geq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{*}(x^{\ast})-\varepsilon.$
Since $\varepsilon>0$ was arbitrary, we deduce that
$f^{\ast}(x^{\ast})\;\geq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{*}(x^{\ast}).$
On the other hand, we always have $f^{\ast}(x^{\ast})\leq{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f_{n}^{*}(x^{\ast})$ (see Lemma 6.1). Combining both inequalities completes the proof.
[…]
7. Proof of Corollary 4.6
Proof. Assume that ${\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}f^{\ast}_{n}=f^{\ast}$. Then
${\Gamma(w^{\ast})\text{-}\displaystyle\liminf_{n}}f^{\ast}_{n}\;\geq\;f^{\ast}\;\geq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f^{\ast}_{n}.$
Taking conjugates and using that ${\rm dom}\kern 1.19995pt({\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f^{\ast}_{n})={\rm dom}\kern 1.19995pt(f^{\ast})\neq\emptyset$, we can apply Theorem 4.5 to obtain
$\Gamma\text{-}\displaystyle\limsup_{n}f_{n}\;\leq\;f\;\leq\;({\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f^{\ast}_{n})^{\ast}.$
Since it is always true that ${\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f^{\ast}_{n}\geq(\Gamma\text{-}\displaystyle\liminf_{n}f_{n})^{\ast}$ (see Lemma 6.1) and also $(\Gamma\text{-}\displaystyle\liminf_{n}f_{n})^{\ast\ast}\leq\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$, we conclude that
$\Gamma\text{-}\displaystyle\limsup_{n}f_{n}\;\leq\;f\;\leq\;\Gamma\text{-}\displaystyle\liminf_{n}f_{n}.$
Therefore,
$f^{\ast}={\Gamma(w^{\ast})\text{-}\displaystyle\lim_{n}}f^{\ast}_{n}\quad\Longrightarrow\quad f=\Gamma\text{-}\displaystyle\lim_{n}f_{n}.$
Now we prove the reverse implication under the assumption that $X$ is a WCG Banach space and $(f^{\ast}_{n})$ satisfies (4.1). Suppose that $f=\Gamma\text{-}\displaystyle\lim_{n}f_{n}$. Then
$\Gamma\text{-}\displaystyle\limsup_{n}f_{n}\;\leq\;f\;\leq\;\Gamma\text{-}\displaystyle\liminf_{n}f_{n}.$
Taking conjugates and applying Theorem 4.1 (since ${\rm dom}\kern 1.19995pt(\Gamma\text{-}\displaystyle\limsup_{n}f_{n})={\rm dom}\kern 1.19995pt(f)\neq\emptyset$), together with Theorem 4.5 (since ${\rm dom}\kern 1.19995pt({\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f^{\ast}_{n})={\rm dom}\kern 1.19995pt(f^{\ast})\neq\emptyset$), we obtain
$({\Gamma(w^{\ast})\text{-}\displaystyle\liminf_{n}}f^{\ast}_{n})^{\ast\ast}\;\geq\;f^{\ast}\;\geq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f^{\ast}_{n}.$
Since $({\Gamma(w^{\ast})\text{-}\displaystyle\liminf_{n}}f^{\ast}_{n})^{\ast\ast}\leq{\Gamma(w^{\ast})\text{-}\displaystyle\liminf_{n}}f^{\ast}_{n}$, we deduce that
${\Gamma(w^{\ast})\text{-}\displaystyle\liminf_{n}}f^{\ast}_{n}\;\geq\;f^{\ast}\;\geq\;{\Gamma(w^{\ast})\text{-}\displaystyle\limsup_{n}}f^{\ast}_{n},$
[…]
8. Proof of Theorem 4.7
In this section we prove Theorem 4.7. Before establishing the implications $(a)\Rightarrow(b)\Rightarrow(c)$, let us first observe that the equivalence of the three items, under the additional equicoercivity assumption (4.1) for the sequence $(f_{n}^{\ast})$, follows directly from Corollary 4.6, which already shows the equivalence between $(a)$ and $(c)$.
Hence, in the remainder of this section we divide the proof of the implications into two subsections.
8.1. Proof of Theorem 4.7: $(a)\Longrightarrow(b)$
Proof. Let $(x,y^{\ast})\in\partial f$, that is,
$f(x)+f^{\ast}(y^{\ast})-\langle y^{\ast},x\rangle=0.$
By hypothesis, there exists a sequence $(z_{n}^{\ast})_{n\in\mathbb{N}}\subset X^{\ast}$ such that $z_{n}^{\ast}\overset{\ast}{\rightharpoonup}y^{\ast}$ and $f^{\ast}_{n}(z_{n}^{\ast})\to f^{\ast}(y^{\ast})$. By Corollary 4.6, there exists a sequence $(u_{n})_{n\in\mathbb{N}}\subset X$ such that $u_{n}\to x$ and $f_{n}(u_{n})\to f(x)$.
Define
$\varepsilon_{n}:=f_{n}(u_{n})+f^{\ast}_{n}(z_{n}^{\ast})-\langle z_{n}^{\ast},u_{n}\rangle>0,$
which is a sequence of positive numbers converging to 0.
By the Brøndsted–Rockafellar theorem (see, e.g.,  [30] or [28, Theorem 4.3.2]), for every $n\in\mathbb{N}$ one can find $(x_{n},y_{n}^{\ast})\in\partial f_{n}$ such that
$\left\lVert z_{n}^{\ast}-y_{n}^{\ast}\right\rVert\leq\sqrt{\varepsilon_{n}}\quad\text{and}\quad\left\lVert x_{n}-u_{n}\right\rVert\leq\sqrt{\varepsilon_{n}}.$
It follows that $x_{n}\to x$ and $y_{n}^{\ast}\overset{\ast}{\rightharpoonup}y^{\ast}$. Since $(x_{n},y_{n}^{\ast})\in\partial f_{n}$, we conclude that
$\partial f_{n}\xrightarrow{G(\|\cdot\|\times w^{\ast})}\partial f.$
[…]
8.2. Proof of Theorem 4.7: $(b)\Longrightarrow(c)$
In order to prove that $(b)\Longrightarrow(c)$ in Theorem 4.7, we will recall the following integration Lemma given in [32, Lemma 2.3]. For that purpose let us introduce the following notation. For a function $\varphi\colon X\to\mathbb{R}\cup\{-\infty,\infty\}$ (not necessarily convex), the Fenchel subdifferential of $\varphi$ at a point $x\in X$ is defined as
$\partial^{\text{Fen}}\varphi(x)=\{x^{*}\in X^{*}:\langle x^{*},u-x\rangle+\varphi(x)\leq\varphi(u),\,\forall u\in X\},$
if $x\in\operatorname{dom}\varphi:=\{x\in X:\varphi(x)<\infty\}$. Otherwise, if $x\notin\operatorname{dom}\varphi$, we set:
$\partial^{\text{Fen}}\varphi(x)=\emptyset.$
Lemma 8.1.
Let $g\colon X\to\mathbb{R}\cup\{\infty\}$ be a lower semicontinuous function and $f:X\to\mathbb{R}\cup\{\infty\}$ be a proper lower semicontinuous convex function. If:
$\partial f\subset\partial^{\text{Fen}}g,$
then $f=g+c$ , for some $c\in\mathbb{R}$.
The following technical lemma will be useful in the sequel. Since its proof is only a slight modification of [32, Lemma 2.4], we omit it here.
Lemma 8.2.
Let $f_{n}\colon X\to\mathbb{R}\cup\{\infty\}$, $n\in\mathbb{N}$, be proper lower semicontinuous convex functions satisfying that $\Gamma\text{-}\displaystyle\liminf_{n}f_{n}>$-$\infty$ and there exists a strongly convergent sequence $(a_{n})_{n\in\mathbb{N}}$ for which
$\limsup_{n}f_{n}(a_{n})<\infty.$
Then, for any sequence $(x_{n},x_{n}^{*})$ with $(x_{n},x_{n}^{*})\in\partial f_{n}$ and $x_{n}\to x$, $x^{\ast}_{n}\overset{\ast}{\rightharpoonup}x^{\ast}$, one has:
  1. (a)
    $\Gamma\text{-}\displaystyle\liminf_{n}f_{n}(x)=\displaystyle\liminf_{n}f_{n}(x_{n})$ and $\Gamma\text{-}\displaystyle\limsup_{n}f_{n}(x)=\displaystyle\limsup_{n}f_{n}(x_{n})$.
  2. (b)
[…]
Furthermore, we have:
$\operatorname{Li}_{\left\lVert\cdot\right\rVert\times w^{\ast}}\partial f_{n}\subset\partial^{\text{Fen}}(\Gamma\text{-}\displaystyle\liminf_{n}f_{n})\cap\partial(\Gamma\text{-}\displaystyle\limsup_{n}f_{n}).$
Now we prove $(b)\Longrightarrow(c)$ from Theorem 4.7. Proof. Indeed, let us assume that (b) holds, i.e.,
$\partial f_{n}\xrightarrow{G(\left\lVert\cdot\right\rVert\times w^{\ast})}\partial f$
and the the normalization condition (N.C.) holds, that is, there exists $(a,a^{*})\in\partial f$ and a sequence $(a_{n},a_{n}^{*})\in\partial f_{n}$ such that $a_{n}\to a,\;a_{n}^{*}\overset{\ast}{\rightharpoonup}a^{*}\text{ and }f_{n}(a_{n})\to f(a).$
We need to show that $f=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}$ and $f=\Gamma\text{-}\displaystyle\limsup_{n}f_{n}$.
The N.C. ensures that the assumptions of Lemma 8.2 are satisfied. Hence, using Lemma 8.2, we deduce that:
$\partial f\subset\operatorname{Li}_{\left\lVert\cdot\right\rVert\times w^{\ast}}\partial f_{n}\subset\partial^{\text{Fen}}(\Gamma\text{-}\displaystyle\liminf_{n}f_{n})\cap\partial(\Gamma\text{-}\displaystyle\limsup_{n}f_{n}).$
From Lemma 8.1, we obtain that:
$f=\Gamma\text{-}\displaystyle\liminf_{n}f_{n}+C_{1}=\Gamma\text{-}\displaystyle\limsup_{n}f_{n}+C_{2},$
where $C_{1}$ and $C_{2}$ are constant real numbers. In addition, Lemma 8.2 the N.C. imply that:
$\Gamma\text{-}\displaystyle\liminf_{n}f_{n}(a)=f(a)=\Gamma\text{-}\displaystyle\limsup_{n}f_{n}(a),$
for some $a\in\operatorname{dom}(f)$. Hence, $C_{1}=C_{2}=0$ and the proof is finished.
[…]